# Mechanically concealed holes

### v1

@@ Abstract
For a detailed overview on the parameter values and protocol of our MD simulations, we refer to the supplemental material [34], which includes the additional Refs. 35, 36, 37, 38, 39, 40, 41. We here just summarize that strongly cohesive, two-dimensional, planar solids consisting of 46200 atoms that interact via a truncated and force-shifted Lennard-Jones (LJ) potential were simulated under periodic boundary conditions. Very low temperatures ($T$ = 0.6 K) were imposed throughout to allow for comparison with the continuum analogs described above. The characteristic size of the atoms was set to $\sigma=1$ Å and the radius of the holes carved out from the solid to $a=20$ Å. Pristine hexagonal solids, hexagonal solids with a hole but no shell ($b/a=1$), and hexagonal solids containing shelled holes ($b/a>1$) were prepared. We here use the bulk modulus to compare their mechanical behavior and test the effective mechanical shielding of the hole from isotropic compression. To introduce stiff shells, we set the minimum of the LJ potential deeper for shell atoms than for atoms belonging to the surrounding elastic solid. At long wavelengths the LJ parameters for a two-dimensional hexagonal lattice can be related to the resulting effective shear modulus [42, 43]. We rely on this relation to define and obtain the ratio between our elastic moduli $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ [34]. Our MD simulations were performed using LAMMPS [44, 45].
@@
Quite remarkably, as shown in Fig. 4, results from atomistic simulations match well the results from analytical continuum elasticity theory over a wide range of ratios $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ for the considered $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ =0.34=0.34. This suggests that the picture provided by continuum elasticity theory remains valid even down to atomistic scales for the considered discrete systems of particles interacting solely via LJ pair potentials. As we gradually reduce the stiffness of the shell surrounding the hole in terms of the ratio $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ in Fig. 4, theory and simulations follow the same overall trend. Yet, quantitative differences appear in the necessary thickness of the shell around the hole that is necessary for effective concealment, as measured by $b/a$. One possible source for these deviations may be found in a progressively pronounced inhomogeneity on the atomic scale, in relation to the finite size of the simulation box. Indeed, the number of atoms within the shell, interacting via stronger interatomic forces compared to the background solid, grows from 0.87 $\%$ to 7.26 $\%$ of the total number of atoms in the solid with lower ratios $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ [34]. Elevated thickness of the shell is necessary to conceal the hole with softer shell materials (lower $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$). Simultaneously, for fixed $a$ (here $20$ Å), elevated $b/a$ implies an increasing outer perimeter of the shell. This implies a larger number of particles interacting inhomogeneously, that is, with different LJ parameters concerning their interaction partners [34]. Conversely, for stiffer and thus thinner shells (higher $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$), the results from continuum theory and atomistic simulations are almost identical.
@@
We remark that further quantitative agreement between analytical results and those from MD simulations can be achieved by reducing the ratio $a/L$ between the radius of the hole and the size of the simulated system. A corresponding demonstration for $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$ is shown in Ref. 34. Yet, generally, if we reduce the size $a$ of the hole, the deviation in bulk elastic modulus is not that large already from the start. In Fig. 3(a) and (b) the red and black curves would be very close to each other, and so would be the blue curve for concealment. Finally, concerning actual parameter values of real materials, a hollow Al shell surrounded by a background solid of epoxy resin would have $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ within the range of 0.3–0.35, and $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}\approx 10$.
@@
Conclusions. Summarizing, we considered a cylindrical hole that is surrounded by a stiffer cylindrical shell in a block of elastic material. The material is loaded uniformly under plane-strain conditions. Our scope was to identify for given material parameters a thickness of the shell that mechanically conceals the presence of the hole on the macroscopic level. In other words, the material as a whole, including such masked holes, behaves mechanically in the same way as if the holes were absent. Such a situation allows reduction in weight of components during light-weight construction, maintaining their overall shape and mechanical properties.
@@
Indeed, a corresponding expression for the thickness of the shell around the hole was identified and derived as a function of the mechanical material parameters and the radius of the concealed hole. For this purpose, we assumed continuous, linearly elastic, homogeneous, isotropic materials under plane-strain conditions. The concentrations of the holes were low enough so that mechanical interactions between the holes can be neglected. One step further, we extended this possibility of mechanical concealment from a macroscopic continuum perspective down to discrete atomistic scales using molecular dynamics simulations. A purely atomistic, planar solid was considered. Consistent results were found. In principle, they validate the predictions of continuum elasticity theory down to the atomistic level. Gradual deviations in our finite-sized atomistic systems emerge at lower shell stiffnesses, which necessitate thicker shells. They imply a larger fraction of the stiffer shell component of the overall system, reducing the relative fraction of the surrounding elastic solid, and pronouncing the role of inhomogeneities.
@@
Future extensions of this work are manifold, important, and obvious. An upcoming investigation shall address more general types of loading, beyond uniform. Besides plane-strain, also plane-stress and three-dimensional situations shall be considered in future works. Elevated concentrations of holes that make them mutually interact through their induced deformations shall be evaluated. More complex considerations include mechanically anisotropic materials such as crystalline solids. Altogether, we wish to support the further development of quantitative measures in the context of light-weight materials design.
@@
Acknowledgments. We thank the European Union (EFRE) and the State Saxony Anhalt for support of this work through project no. ZS/2024/02/184030. Moreover, we thank Lukas Fischer, Johannes Menning, and Thomas Wallmersperger for stimulating discussions.

### v2

@@ III.4 MD results: shell thickness for concealment
Figure 6: Atomistic distributions of instantaneous virial shear stress during isotropic compression at an areal strain of 0.3 $\%$. We consider (a) a pristine elastic solid without any hole, (b) an elastic solid with an unshielded hole ($b/a$ = 1), (c) an elastic solid with a shielded hole of parameters $b/a=1.173$, $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=10$, and (d) an elastic solid with a shielded hole and parameters $b/a=2.018$, $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$. The outer perimeter of the shielding shell of radius $b$ is shown via a black, dashed circle in (c) and (d).
Moreover, we investigate the effect of mechanical shielding of holes on the stress distribution of the solid on the atomistic level. We recall that the shielding strategy relies on tuning the ratio $b/a$, so that the effective bulk modulus matches that of the pristine solid. Thereby, we account for the stress components $\sigma_{xx}$ and $\sigma_{yy}$. Ideally, macroscopic isotropic compression does not intend to impose global shear stresses. Still, fluctuations on the atomic level persist and could give rise to microscopic, instantaneous virial shear stresses around the hole. This naturally raises the more specific, subsequent question about whether the cloak mitigates these instantaneous virial shear stresses during the compressive response at the atomistic level.
Figure 6 represents the distribution of virial shear stress on the atoms during isotropic compression for pristine, unshielded as well as for concealed holes. The areal strain is set to 0.3 $\%$ to remain within the elastic limit. As a measure for the atomic shear stress, we follow the expression Thompson, Plimpton, and Mattson (2009)
$\tilde{A}\sigma_{xy}^{\text{at}}=\sum_{i=1}^{N}\left[{}-m_{i}v_{i,x}v_{i,y}-\sum_{j>i}r_{ij,x}F_{ij,y}\right].$ (43)
Here, $\tilde{A}$ is a typical area of the order of the area of one atom. $m_{i}$ is the mass of particle $i$. $x$- and $y$-components of the velocity of particle $i$ are defined as $v_{i,x}$ and $v_{i,y}$, respectively. $r_{ij,x}$ denotes the $x$-component of the distance and $F_{ij,y}$ describes the $y$-component of force between particles $i$ and $j$. Since the kinetic term is typically very small, we neglect it in the calculation of the shear stress.
The pristine solid without the hole in Fig. 6(a) shows a rather scattered distribution of local stress. Contrarily, instantaneous shear stresses are observed to build up and accumulate around an unshielded hole, see Fig. 6(b). They reach into the surrounding solid to a notable distance. Upon cloaking the hole, that is, inserting the stiffer shell, we observe that the desired $b/a$ for concealment effectively confines the instantaneous virial shear stresses around the hole to the region within the shell thickness, see Fig. 6(c,d). The background solid remains in these cases with similar scattered distributions of virial shear stress as for the pristine solid without the hole in Fig. 6(a).
IV Concealment ratio: comparing atomistic and continuum results
Quite remarkably, as shown in Fig. 7, results from atomistic simulations match well the results from analytical continuum elasticity theory over a wide range of ratios $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ for the considered $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ =0.34=0.34. This suggests that the picture provided by continuum elasticity theory remains valid even down to atomistic scales for the considered discrete systems of particles interacting solely via LJ pair potentials. As we gradually reduce the stiffness of the shell surrounding the hole in terms of the ratio $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ in Fig. 7, theory and simulations follow the same overall trend. Yet, quantitative differences appear in the necessary thickness of the shell around the hole that is necessary for effective concealment, as measured by $b/a$.
@@ IV Concealment ratio: comparing atomistic and continuum results
One possible source for these deviations may be found in a progressively pronounced inhomogeneity on the atomic scale, in relation to the finite size of the simulation box. Indeed, the number of atoms within the shell, interacting via stronger interatomic forces compared to the background solid, grows from 0.87 $\%$ to 7.26 $\%$ of the total number of atoms in the solid with lower ratios $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$, see Fig. 8(a). Elevated thickness of the shell is necessary to conceal the hole with softer shell materials (lower $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$). Simultaneously, for fixed $a$ (here $20$ Å), as shown in Fig. 8(b), elevated $b/a$ implies an increasing outer perimeter of the shell. This implies a larger number of particles interacting inhomogeneously, that is, with different LJ parameters concerning their interaction partners . Conversely, for stiffer and thus thinner shells (higher $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$), the results from continuum theory and atomistic simulations are almost identical.
Figure 8: (a) Number of atoms within the shell (Nshell) relative to the total number of atoms in the system (N) as a function of the ratio of shear moduli $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$ when the presence of the hole is effectively concealed. (b) For the same systems variation of the required outer perimeter of the shell as a function of $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. Dashed black lines are guides to the eye.
We remark that further quantitative agreement between analytical results and those from MD simulations can be achieved by reducing the ratio $a/L$ between the radius of the hole and the size of the simulated system. A corresponding demonstration for $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$ is shown in Fig. 9, where we observe that this difference gradually decreases as we decrease $a/L$ from 0.089 to 0.009. We remark that, for this purpose, in our simulations we stepwise reduced the radius of the hole $a$ as indicated in the legend.
@@
Yet, generally, if we reduce the size $a$ of the hole, the deviation in bulk elastic modulus is not that large already from the start. In Fig. 5(a) and (b) the red and black curves would be very close to each other, and so would be the blue curve for concealment. Finally, concerning actual parameter values of real materials, a hollow Al shell surrounded by a background solid of epoxy resin would have $\nu^{\mathrm{i}}$ $\approx$ $\nu^{\mathrm{o}}$ within the range of 0.3–0.35, and $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}\approx 10$.
Figure 9: Ratio $b/a$ between outer radius of the shell $b$ and the radius of the hole $a$ to mechanically conceal the hole under isotropic compression. With decreasing $a$, its ratio with respect to the size of the MD system $L$ decreases as well. We chose a specific, low value of the stiffness ratio $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=2$. The blue line marks the result from continuum elasticity theory.
V Conclusions
@@
Summarizing, we considered a cylindrical hole that is surrounded by a stiffer cylindrical shell in a block of elastic material. The material is loaded uniformly under plane-strain conditions. Our scope was to identify for given material parameters a thickness of the shell that mechanically conceals the presence of the hole on the macroscopic level. In other words, the material as a whole, including such masked holes, behaves mechanically in the same way as if the holes were absent. Such a situation allows reduction in weight of components during light-weight construction, maintaining their overall shape and mechanical properties.
@@
Indeed, a corresponding expression for the thickness of the shell around the hole was identified and derived as a function of the mechanical material parameters and the radius of the concealed hole. For this purpose, we assumed continuous, linearly elastic, homogeneous, isotropic materials under plane-strain conditions. The concentrations of the holes were low enough so that mechanical interactions between the holes can be neglected. One step further, we extended this possibility of mechanical concealment from a macroscopic continuum perspective down to discrete atomistic scales using molecular dynamics simulations. A purely atomistic, planar model solid was considered. Although gradual deviations in our finite-sized atomistic systems emerge at lower shell stiffnesses, which necessitate thicker shells when particulate inhomogeneities play a larger role, overall consistent results were found. In principle, they validate the concept introduced via continuum elasticity theory down to the atomistic level.
As a central step, we thus resolve the concept predicted by continuum elasticity theory on a discrete particulate scale by molecular dynamics simulations. While the continuum elasticity solution renders the problem tractable with only little effort in optimization when adjusting the shell thickness, molecular dynamics simulations motivate broad opportunities for microscopic manipulations that may lead to innovative cloaking techniques.
@@
Future extensions of this work are manifold, important, and obvious. An upcoming investigation shall address more general types of loading, beyond uniform. Besides plane-strain, also plane-stress and three-dimensional situations shall be considered in future works. Elevated concentrations of holes that make them mutually interact through their induced deformations shall be evaluated. More complex considerations include mechanically anisotropic materials. Altogether, we wish to support the further development of quantitative measures in the context of light-weight materials design.
Acknowledgements.
@@
We thank the European Union (EFRE) and the State Saxony Anhalt for support of this work through project no. ZS/2024/02/184030. Moreover, we thank Lukas Fischer, Johannes Menning, and Thomas Wallmersperger for stimulating discussions.
Data Availability Statement
The generated molecular dynamics simulation data, used to derive the results presented herein, including all the figures, are openly available in the repository Zenodo and can be found at https://doi.org/10.5281/zenodo.17957498.
Appendix A Lennard-Jones “toy model” for a strongly cohesive solid with a hole
Atomistic modeling of holes inside solids is challenging, given that the holes could deform, distort in search of the thermodynamically equilibrated state. Therefore, such geometries are obscure in the literature on MD simulations. Numerical investigations of mechanical concealment or cloaking often consider specific materials, such as graphene, using the finite-element approach (Bückmann et al., 2015). Furthermore, covalently bonded materials necessitate the usage of many-body interatomic potentials. They can be stabilized with voids in MD simulations, but could invoke anisotropy due to directional bonding and consequently having interatomic potentials with angular dependence.
Choosing specific combinations of materials in atomistic simulations in our specific geometry of a hole, shielded by an elastic shell, embedded in a surrounding elastic solid, limits, for each realization, the comparison between continuum theory and atomistic simulations to a specific ratio of the material parameters $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}$. To this end, we choose to model a physically traceable system that is simple enough to retain isotropy (potential without angular dependence), yet complex enough to allow mechanical concealment using a shell of variable stiffness around a hole.
We employ a two-dimensional Lennard-Jones (LJ) solid to mimic a strongly cohesive solid. In Eqs. (35) and (36), we choose $\sigma$ = 1 Å for both types of atoms, part of the shell or the surrounding solid. We modify the energy scale (depth of the energy well), mimicking a “toy model” for a strongly cohesive solid with $\epsilon^{\mathrm{o}}=1$ eV and $2\epsilon^{\mathrm{o}}\leq\epsilon^{\mathrm{i}}\leq 15\epsilon^{\mathrm{o}}$, where $\epsilon^{\mathrm{i,o}}$ quantifies the stiffness between the atoms within the shell (i) and within the background solid (o), see Fig. 10. Similar approaches to control the interparticle stiffness using a single parameter in a pairwise potential are found in the literature, albeit in a different context of modeling binary glass formers with a slightly different shape of the pair potential (Lerner, 2022).
Figure 10: Truncated and shifted Lennard-Jones potential, see Eqs. (35) and (36). The curves illustrate realizations for different values of the parameter $\epsilon$ quantifying the depth of the energy well. Apart from that, the length parameter $\sigma=1$, equilibrium distance $r_{0}=2^{1/6}\sigma$ (gray dashed line and bold dots), and cutoff distance $r_{c}$ (red dotted line) are identical in all cases. In our realization, different stiffness parameters $\epsilon$ are used, depending on whether the interacting atoms are parts of the shell around the hole or of the surrounding elastic solid.
Indeed, stability of a LJ solid in two dimensions with a void is unattainable using the standard cohesive energy values of rare-gas solids ($\epsilon\sim 10^{-2}$ eV for Ar). Therefore, we use deeper energy wells, which results in very high elastic moduli of the order of TPa. We remark, that monolayer graphene, having a honeycomb hexagonal structure, possesses similarly high elastic moduli of the order of TPa. Our two-dimensional LJ model of a strongly cohesive solid is even stiffer. In our case, we study a hexagonal, six-neighbor structure, as opposed to the three-neighbor structure in monolayer graphene.
Overall, we thus remark that the parameters for our LJ solid were adjusted to maintain a simple, basic description as a proof of concept on the atomic scale. Although, the considered parameters do not reflect a realistic solid at these elevated values for the elastic moduli, we note that the LJ energy parameters $\epsilon$ for metals (Jacobson and Thompson, 2022a, b; Filippova, Kunavin, and Pugachev, 2015) and even for two-dimensional materials (Momeni et al., 2020) often correspond to such high values ranging mostly within $0.1$ eV $\epsilon<1.5$ eV.
As an example, we include a relative comparison between $\epsilon^{\mathrm{o}}=0.01$ and 1 eV in sustaining a shielded hole under NPT equilibration at $P=2\times 10^{5}$ bar in Fig. 11. Reducing $\epsilon^{\mathrm{o}}$ to 0.1 eV can sustain the hole with slightly lower bulk modulus. Yet, we chose to use $\epsilon^{\mathrm{o}}=1$ eV and $\sigma$ = 1 Å so that the problem can be conveniently expressed in reduced units. Similarly, larger values of $\epsilon$ were used to simulate two-dimensional LJ solids with periodically arranged square holes in the context of stress analysis in inhomogeneous materials (Rigelesaiyin et al., 2018).
Figure 11: Undistorted and distorted hole under NPT equilibration at $P=2\times 10^{5}$ bar. We display parts of planar solids with a hole for an LJ energy parameter of the background solid (a) $\epsilon^{\mathrm{o}}=1$ eV and (b) $\epsilon^{\mathrm{o}}=0.01$ eV. $\epsilon^{\mathrm{i}}/\epsilon^{\mathrm{o}}=10$ in both cases.
Figure 12: Radial distribution function g(r) for (a) a pristine planar LJ solid, (b) a solid with an unshielded hole ($b/a=1$), and (c) a solid with a shielded hole with $b/a=1.173$ and $\mu^{\mathrm{i}}/\mu^{\mathrm{o}}=10$. Insets show parts of the corresponding atomistic systems. Lattice parameters agree with the corresponding minima of the truncated and shifted LJ energies ($r_{0}=1.12\sigma$). The second-neighbor distances are consistent with hexagonal symmetry.
Consistently with the chosen value of $\sigma$, all atomistic systems simulated in this work possess a constant lattice parameter $r_{0}=2^{1/6}\sigma$. This can be inferred from the radial distribution function g(r) computed for different systems with and without holes, see Fig. 12.